What are coplanar and triple product of vector and their relationship in mathematics fotm???????

Grade:11

3 Answers

Aditya Gupta
5 years ago
3 vectors are coplanar if and only if ×c)scalar triple product of vectors is defined as a.(b×c)=0a.(b
Aditya Gupta
5 years ago
due to technical error the above ans is not clear.
scalar triple product of three vectors is defined as:
a.(b×c)
they are coplanar if and only if
a.(b×c)=0
Bhagyashri verma
5 years ago
Hi,
 
There are two triple products
  1. Scalar triple product
  2. vector triple product
Scalar triple product: a.(b X c) 
Vector triple product : \mathbf{a}\times (\mathbf{b}\times \mathbf{c}) = \mathbf{b}(\mathbf{a}\cdot\mathbf{c}) - \mathbf{c}(\mathbf{a}\cdot\mathbf{b}).
 
The vectors are coplanar if scalar triple product is zero